OpenStudy (anonymous):

A right triangle has base (x - 5) units and height (4x - 10) units. Part A: What is the square of the length of the hypotenuse of the triangle?

Vocaloid (vocaloid):

"square of the length of the hypotenunse" = (base)^2 + (height)^2

OpenStudy (anonymous):

(x-5)^2 + (4x -10)^2

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

correct

OpenStudy (anonymous):

Solve what you've written out.

Vocaloid (vocaloid):

oh, well, I guess you could try expanding it (x-5)^2 = (x-5)(x-5) = ? (4x-10)^2 = (4x-10)(4x-10) = ?

OpenStudy (anonymous):

x^2-10x+25 x^2 -20x+100

Vocaloid (vocaloid):

first one looks good check the second one again

OpenStudy (anonymous):

oh 4x sorry

Vocaloid (vocaloid):

nope, not quite... (4x)*(4x) = ?

OpenStudy (anonymous):

no i did x instead of 4x

Vocaloid (vocaloid):

let's start again (4x-10)^2 = (4x-10)(4x-10) = ?

Vocaloid (vocaloid):

remember foil? first we multiply the first terms in each parentheses (4x)*(4x) = ?

OpenStudy (anonymous):

16x^2-80x+100

Vocaloid (vocaloid):

yup! that looks much better now we have: x^2-10x+25 16x^2-80x+100

OpenStudy (anonymous):

Okay

Vocaloid (vocaloid):

now we add everything together combine like terms: x^2 + 16x^2 = ?

OpenStudy (anonymous):

16x^2-90x+125 i forgot if you add the exponents or not

OpenStudy (anonymous):

17x^2-90x+125 sorry

OpenStudy (anonymous):

@Vocaloid is that it?

Vocaloid (vocaloid):

yup! good job

OpenStudy (anonymous):

That doesnt seem like the right answer to the question but okay